# Modelling and Classifying the Components of a Literature Review

### v1

@@ 5.2 Fine-tuning Experiments
6 Additional Analysis
This section provides a detailed analysis of the models’ performance from multiple perspectives. In particular, Section 6.1 investigates the common errors made by the top-performing LLMs. Section 6.2 explores the effects of various optimization techniques, while Section 6.3 assesses the contribution of augmented data to this task.
@@ 6.1 Error Analysis through Confusion Matrices
Sentences annotated as Limitation in the gold standard were most often misclassified as Research Gap or Result. A detailed analysis of these cases suggests that this happens because Limitation are often phrased in a way that implies a broader lack of knowledge, which can be interpreted as a Research Gap. Moreover, some Limitation explicitly refer to the results, which can lead to their misclassification as belonging to the Result category.
Similarly, sentences annotated as Description were most frequently misclassified as Overall or Extension. Our manual analysis indicates that some Description lack sufficient context, making them appear to describe the overall status of the topic and thus leading to their assignment to the Overall category. In other cases, Description outline the study’s design in a way that resembles an expansion of previous methodology, aligning with the definition of the Extension category.
@@ 6.2 Comparing Optimisation Techniques
In conclusion, the evidence is insufficient to definitively establish the superiority of one optimization technique over the other for small and medium models. However, for large decoder models, LoRA produced more favorable results.
@@
6.3 Assessing the Efficacy of Syntetic Data
Semi-synthetic training data produced by LLMs have proven to be very effective, but their performance is inconsistent across different classification tasks [97]. One of our goals was to determine whether semi-synthetic data within Sci-Sentence, generated by creating alternative versions of manually classified sentences from the original data, could improve performance.
To this end, we trained each model on both the augmented and the original training data. As for the optimization techniques, in the previous sections we always considered the best-performing model between the two configurations. Table 9 compares the F1 scores of the models trained on the augmented data and on the original data. The Performance Gain row refers to the difference between the F1 score obtained with the augmented data and the F1 score obtained with the original data. A positive value, therefore, indicates an improvement in performance due to the augmented data.
@@ 6.3 Assessing the Efficacy of Syntetic Data
Encoder-decoder models also tend to benefit from augmented data, especially in their smaller variants, but not to the same extent as encoder models. In sum, our results of semi-synthetic data increasing the performance of encoders or encoder-decoder architectures align with existing literature [114, 115, 116].
For decoder models, the performance gains are more variable. This observation is consistent with the literature: while some studies report a positive effect [117, 118, 119, 120], others find little or no benefit [121, 122, 123, 124, 125]. The greatest benefits are observed in the smaller variants: three out of seven (namely TinyLlama, SmolLM2, and Gemma2-2B) show notable improvements, with TinyLlama achieving a substantial F1-score increase of over 30%. Among the medium-sized models, only two exhibit improvements. It is interesting to note that one of them is Nemotron-8B, which is also the best-performing model in this category.
@@ 7 Conclusion
As future work, we plan to advance our research on multiple fronts. First, we aim to extend the classifier from single-label to multi-label in order to better capture the multifaceted nature of complex sentences that may pertain to multiple categories. Second, we plan to investigate how to capture more elusive categories, such as critiques and interpretations of a piece of literature, which are crucial for incorporating critical evaluations, nuanced perspectives, and broader contextual understanding into our representation. Although existing work, such as Khoo et al. [23], identifies an author’s ‘interpretation’ category, we argue that a more fine-grained approach is necessary to effectively support the automatic generation of literature reviews, instead of aggregating all interpretations into a single category. Finally, we intend to develop a novel framework for generating automatic literature reviews that integrates the framework presented in this paper and moves beyond simple multi-document summarisation towards producing high-quality, in-depth analyses of the literature.
@@ E.1.1 Full Models
Llama 2 Chat 70B888 Llama 2 Chat 70B - (https://llama.meta.com/llama2/) (shortened as llama2) has been trained using a dataset comprising 2 trillion tokens derived from publicly accessible online sources. It has 70 billion parameters and a context length of 4,096 tokens.
Llama 3 8B Instruct999 Llama 3 8b Instruct - (https://ai.meta.com/blog/meta-llama-3/) (shortened as llama3 8B full) is an auto-regressive language model that employs an optimised transformer architecture with 8 billion parameters. It possesses a context length of 8,000 tokens and has been trained on a dataset consisting of 15 trillion tokens. The training process involved supervised fine-tuning (SFT) in conjunction with reinforcement learning from human feedback (RLHF) to align the model with human preferences for utility and safety. To enhance inference scalability, the model incorporates Grouped-Query Attention (GQA).
@@
Mistral 7b Instruct [126] (shortened as mistral) is a 7 billion parameter model with a context length of 32,000 tokens. It incorporates architectural innovations such as Sliding Window Attention mechanism, GQA, and Byte-fallback Byte Pair Encoding (BPE) tokenizer. The first architectural innovation, accommodates a context length of 8,000 tokens and features a fixed cache size, which theoretically enables it to process up to 128,000 tokens. The second innovation, enhances inference speed while reducing cache size. While the third innovation, ensures reliable character recognition without the need for out-of-vocabulary tokens.
Mixtral 8X7b Instruct [127] (shortened as mixtral) consists of 46.7 billion parameters and is capable of processing a context length of 32,000 tokens. It is based on a Sparse Mixture of Experts (SMoE) architecture with open weights. This model employs the same innovative architecture as the Mistral 7b Instruct. However, the Sliding Window Attention mechanism restricts it to handling a context length of 8,000 tokens.
Mistral Large111111 Mistral Large - https://mistral.ai/news/mistral-large/ features a context length of 32,000 tokens and employs 123 billion parameters. The model was trained using a heterogeneous dataset that encompassed a substantial amount of code, multilingual information, and content across a broad spectrum of topics.
@@ E.1.2 Quantised Models
Gemma2-2B-Instruct121212 Gemma2-2B-Instruct- https://huggingface.co/google/gemma-2b-it-GGUF(shortened as Gemma) is a quantized version of Google’s Gemma-2b-it language model [128] converted to the GPT-Generated Unified Format (GGUF) file format with 2 billion parameters and context length of 8,192 tokens.
Orca-2-13B131313 Orca-2-13B - https://huggingface.co/TheBloke/Orca-2-13B-GGUF(shortened as Orca) It is a quantized version of Microsoft’s Orca 2 [129], converted to the GGUF format having 13 billion parameters and a context length of 4096 tokens [130]. It was fine-tuned on LLama 2 13B base model.
@@ E.1.3 Proprietary Models
Sonnet 3.0 (Sonnet) and Haiku 3.0 (Haiku) are part of the Claude 3 series, a family of LLMs developed by Anthropic [131]. Both models were trained on a proprietary corpus derived from both public and private sources. They have a context window of 200,000 tokens. Their distinction resides in the fact that Haiku is designed for immediate responses, while Sonnet 3.0 possesses the capability to manage complex tasks thanks to its architectural framework. These models adhered to the Constitutional AI framework, ensuring their alignment with the principles of helpfulness, honesty, and non-harmfulness
GPT-3.5 Turbo (GPT-3.5) and GPT-4 Turbo (GPT-4), both developed by OpenAI, exhibit specific distinctions in their design and capabilities [132]. GPT-3.5 Turbo operates with a context window of 16,000 tokens, whereas GPT-4 Turbo features a larger context window of 128,000 tokens. Furthermore, GPT-4 Turbo is capable of processing both textual and visual inputs, whereas GPT-3.5 Turbo is restricted to textual data exclusively. Additionally, GPT-4 was trained on a larger and more heterogeneous dataset compared to GPT-3.5 Turbo. Despite this, GPT-3.5 Turbo continues to be a practical and economical choice for numerous applications because of its effective combination of performance and efficiency.
@@ E.2 Decoder-Smal FT
The number of parameters in this category ranges from 1 billion to 3.8 billion. However, the context window and the number of training tokens differ across models. For instance, Olmo-1B-Instruct (Olmo-1B) [133], with 1 billion parameters, and TinyLlama-1.1B-Chat-v1.0 (TinyLlama) [134], with 1.1 billion parameters, share the same context window size of 2,048 tokens and were trained on 3 trillion tokens. In contrast, SmolLM2141414 SmolLM2 - https://github.com/huggingface/smollm/blob/main/README.md, which has 1.7 billion parameters, also employs a context window of 2,048 tokens but was trained on a substantially larger dataset of 11 trillion tokens. These models also vary in their training sources: Olmo-1B was trained on subsets of Dolma v1.7 [135], TinyLlama utilized the architecture and tokenizer of Llama 2 [136], and SmolLM2 was trained on a diverse dataset that includes textbooks, web content, code, mathematics, and external data sources.
On the other hand, Llama3.2-3-Instruct (Llama3.2-3B)151515 Llama3.2-3B-Instruct - https://huggingface.co/meta-llama/Llama-3.2-3B-Instruct was trained on a dataset comprising 15 trillion tokens, with a parameter count of 3 billion and a context window extending up to 8,000 tokens. In contrast, Phi3.5-mini-Instruct (Phi3.5) [^16] Phi3.5-mini-Instruct - https://huggingface.co/microsoft/Phi-3.5-mini-instruct was trained on 3.4 trillion tokens, incorporating 3.8 billion parameters and a significantly larger context window of 128,000 tokens. Gemma2-2B-Instruct (Gemma2-2B) shares the same characteristics as those described in the category decoder OSL, differing only in its data source171717 Gemma2-2B-Google - https://huggingface.co/google/gemma-2-2b-it. The only merge model in this category is Arcee-Lite181818 Arcee-Lite - https://huggingface.co/arcee-ai/arcee-lite which have 1.5 billion parameters and was developed using Distilkit191919 Distilkit - https://github.com/arcee-ai/DistillKit. It supports a context window of 32,000 tokens and its distillation source is Phi-3-Medium [137].
@@ E.3 Decoder-Medium FT
For merged models, this category includes Arcee-Spark [^21] Arcee-Spark - https://huggingface.co/arcee-ai/Arcee-Spark and Llama-3.1-SuperNova-Lite (SuperNova-Lite) [^22] SuperNova-Lite - https://huggingface.co/arcee-ai/Llama-3.1-SuperNova-Lite. Arcee-Spark, initialized from the Qwen2-7B-Instruct [138], comprises 7 billion parameters. On the other hand, SuperNova-Lite, with 8 billion parameters, is built upon the Llama-3.1-8B-Instruct [^23] Llama-3.1 - https://ai.meta.com/blog/meta-llama-3-1/ architecture and distilled from the Llama-3.1-405B-Instruct model. Both models have a context window of 128,000 tokens.
@@ E.4 Decoder-Large FT
This category includes models with parameters ranging from 9 billion to 14 billion. In this sense, Mistral-Nemo-Instruct-2407 (Mistral-Nemo)242424 Mistral Nemo - https://mistral.ai/news/mistral-nemo/ has 12 billion parameters and a 128,000 token context window. Its training dataset incorporates a mix of multilingual text, code data, and conversational-style data to ensure high-quality input. Whereas, SuperNova-Medius252525 Mistral Nemo - https://huggingface.co/arcee-ai/SuperNova-Medius is a merged model of 14 billion parameters based on the Qwen2.5-14B-Instruct architecture. It combines knowledge from both the Qwen2.5-72B-Instruct model and the Llama-3.1-405B-Instruct model through distillation. It has a context windows of 131,072 tokens. In contrast, Gemma2-9-Instruct (Gemma2-9B) has the same features as in Gemma, but differs on its source262626 Gemma2-9B-Google - https://huggingface.co/google/gemma-2-9b-it and the number of parameters which are 9 billion. The only proprietary model is GPT-4o-mini-2024-07-18 (GPT-4o-mini)272727 GPT-4o-mini - https://openai.com/index/gpt-4o-mini-advancing-cost-efficient-intelligence/ which is part of the GPT-4o family and is designed to be a cost-efficient, high-performance AI model. It has multimodal capabilities and its context window of 128,000 tokens. It is designed to replace GPT-3.5 Turbo in ChatGPT due to its improved performance and cost-efficiency for various AI applications.

### v2

@@ 5.4 Error Analysis through Confusion Matrices
Sentences annotated as Limitation in the gold standard were most often misclassified as Research Gap or Result.
A detailed analysis of these cases suggests that this occurs because instances of Limitation are often phrased in a way that implies a broader lack of knowledge. This phrasing can lead them to be interpreted as a Research Gap. For example, consider the sentence: “Although the accuracy of current shape segmentation exceeds 90%, some models might still be poorly segmented and thus fail to be accurately landmarked”. The intended meaning is to describe a limitation of the specific model being examined. However, the generalized expression ”current shape segmentation” implies a shortcoming affecting the entire field rather than a precise constraint of the study. As a consequence, such statements may be misclassified as a Research Gap, since they seem to highlight an area where further research or improvement is needed across existing approaches instead of pointing to a limitation of the work at hand.
Moreover, some Limitation sentences explicitly refer to the results, which can lead to their misclassification as belonging to the Result category. For instance, the sentence “Our encoding of traces does allow for uneven data samples, but it is probable that the performance will be reduced when sampling becomes irregular or infrequent” can be misclassified as a Result, because it reports a finding about the model’s ability to handle uneven data. However, it simultaneously conveys a limitation by acknowledging that performance may deteriorate under certain conditions, thus blurring the distinction between reporting results and discussing constraints. This characteristic of sentences belonging to multiple rhetorical types is a specific challenge that we aim to address in future refinement efforts.
On the other hand, sentences annotated as Description were most frequently misclassified as Overall or Extension. Our manual analysis indicates that some Description sentences can be confused with introductory clauses that appear to summarise the general status of a method or topic. This can lead to their incorrect assignment to the Overall category. This issue is exemplified by the sentence “Specifically, as the first inductive kg embedding method, mean [8] learns to represent entities using their neighbors by simply mean-pooling the information of neighboring entity-relation pairs”. The introductory phrase presents the method in a broad and contextual manner, highlighting its novelty or general contribution rather than providing a precise descriptive detail. This framing causes the sentence to resemble a high-level summary, which can lead to its misclassification as Overall. However, the latter part of the sentence (“learns to represent entities using their neighbors…”) in fact specifies how the method operates, which is characteristic of the Description category.
In other cases, Description outline the study’s design in a way that resembles an expansion of previous methodology, aligning with the definition of the Extension category. For example, the sentence “Telm [42] improves on TComplEx by utilizing a linear temporal regularizer and multi-vector embeddings to perform 4th-order tensor factorization of TKGs” might appear to fit the Extension category because it references a previous method. However, it should be classified as Description, since it does not explain how the current study elaborates on or extends the previous one. For a sentence to qualify as Extension, it must detail the nature of the elaboration, for example, by clarifying what limitations are addressed or what specific aspects of the prior study are developed further.
@@ 5.5 Comparing Optimisation Techniques
In conclusion, the evidence is insufficient to definitively establish the superiority of one optimization technique over the other for small and medium models. However, for large decoder models, LoRA produced more favourable results.
@@
5.6 Assessing the Efficacy of Synthetic Data
Semi-synthetic training data produced by LLMs have proven to be very effective, but their performance is inconsistent across different classification tasks (Li et al., 2023b, ). One of our goals was to determine whether semi-synthetic data within Sci-Sentence, generated by creating alternative versions of manually classified sentences from the original data, could improve performance.
To this end, we trained each model on both the augmented and the original training data. As for the optimization techniques, in the previous section we always considered the best-performing model between the two configurations. Table 9 compares the F1 scores of the models trained on the augmented data and on the original data. The Performance Gain row refers to the difference between the F1 score obtained with the augmented data and the F1 score obtained with the original data. A positive value, therefore, indicates an improvement in performance due to the augmented data. It is important to note that Sonnet 3.0, which was used to generate the augmented datasets, was not included among the fine-tuning models in order to avoid potential issues of circularity (Clarke and Dietz, , 2024).
@@ 5.6 Assessing the Efficacy of Synthetic Data
Encoder-decoder models also tend to benefit from augmented data, especially in their smaller variants, but not to the same extent as encoder models. In sum, our results of semi-synthetic data increasing the performance of encoders or encoder-decoder architectures align with existing literature (Fields et al., , 2024; Chae and Davidson, , 2023; Fatemi et al., , 2024).
For decoder models, the performance gains are more variable. This observation is consistent with the literature: while some studies report a positive effect (Singh et al., , 2023; Li et al., 2024b, ; Zhezherau and Yanockin, , 2024; Chen et al., , 2024), others find little or no benefit (Guo et al., , 2023; Zhao et al., 2024b, ; Li et al., 2024a, ; Mecklenburg et al., , 2024; Maheshwari et al., , 2024). The greatest benefits are observed in the smaller variants: three out of seven (namely TinyLlama, SmolLM2, and Gemma2-2B) show notable improvements, with TinyLlama achieving a substantial F1-score increase of over 30%. Among the medium-sized models, only two exhibit improvements. It is interesting to note that one of them is Nemotron-8B, which is also the best-performing model in this category.
@@ 6 Conclusion
As future work, we plan to advance our research on multiple fronts. First, we aim to extend the classifier from single-label to multi-label in order to better capture the multifaceted nature of complex sentences that may pertain to multiple categories. Second, we plan to investigate how to capture more elusive categories, such as critiques and interpretations of a piece of literature, which are crucial for incorporating critical evaluations, nuanced perspectives, and broader contextual understanding into our representation. Although existing work, such as Khoo et al., (2011), identifies an author’s ‘interpretation’ category, we argue that a more fine-grained approach is necessary to effectively support the automatic generation of literature reviews, instead of aggregating all interpretations into a single category. Finally, we intend to develop a novel framework for generating automatic literature reviews that integrates the framework presented in this paper and moves beyond uncritical multi-document summarisation towards producing high-quality, in-depth analyses of the literature.
@@ E.1.1 Full Models
Llama 2 Chat 70B [^8] Llama 2 Chat 70B - (https://llama.meta.com/llama2/) (shortened as llama2) has been trained using a dataset comprising 2 trillion tokens derived from publicly accessible online sources. It has 70 billion parameters and a context length of 4,096 tokens.
Llama 3 8B Instruct [^9] Llama 3 8b Instruct - (https://ai.meta.com/blog/meta-llama-3/) (shortened as llama3 8B full) is an auto-regressive language model that employs an optimised transformer architecture with 8 billion parameters. It possesses a context length of 8,000 tokens and has been trained on a dataset consisting of 15 trillion tokens. The training process involved supervised fine-tuning (SFT) in conjunction with reinforcement learning from human feedback (RLHF) to align the model with human preferences for utility and safety. To enhance inference scalability, the model incorporates Grouped-Query Attention (GQA).
@@
Mistral 7b Instruct (Jiang et al., , 2023) (shortened as mistral) is a 7 billion parameter model with a context length of 32,000 tokens. It incorporates architectural innovations such as Sliding Window Attention mechanism, GQA, and Byte-fallback Byte Pair Encoding (BPE) tokenizer. The first architectural innovation, accommodates a context length of 8,000 tokens and features a fixed cache size, which theoretically enables it to process up to 128,000 tokens. The second innovation, enhances inference speed while reducing cache size. While the third innovation, ensures reliable character recognition without the need for out-of-vocabulary tokens.
Mixtral 8X7b Instruct (Jiang et al., , 2024) (shortened as mixtral) consists of 46.7 billion parameters and is capable of processing a context length of 32,000 tokens. It is based on a Sparse Mixture of Experts (SMoE) architecture with open weights. This model employs the same innovative architecture as the Mistral 7b Instruct. However, the Sliding Window Attention mechanism restricts it to handling a context length of 8,000 tokens.
Mistral Large [^11] Mistral Large - https://mistral.ai/news/mistral-large/ features a context length of 32,000 tokens and employs 123 billion parameters. The model was trained using a heterogeneous dataset that encompassed a substantial amount of code, multilingual information, and content across a broad spectrum of topics.
@@ E.1.2 Quantised Models
Gemma2-2B-Instruct [^12] Gemma2-2B-Instruct- https://huggingface.co/google/gemma-2b-it-GGUF(shortened as Gemma) is a quantized version of Google’s Gemma-2b-it language model (Team et al., , 2024) converted to the GPT-Generated Unified Format (GGUF) file format with 2 billion parameters and context length of 8,192 tokens.
Orca-2-13B [^13] Orca-2-13B - https://huggingface.co/TheBloke/Orca-2-13B-GGUF(shortened as Orca) It is a quantized version of Microsoft’s Orca 2 (Mukherjee et al., , 2023), converted to the GGUF format having 13 billion parameters and a context length of 4096 tokens (Mitra et al., , 2023). It was fine-tuned on LLama 2 13B base model.
@@ E.1.3 Proprietary Models
Sonnet 3.0 (Sonnet) and Haiku 3.0 (Haiku) are part of the Claude 3 series, a family of LLMs developed by Anthropic (Anthropic, , 2024). Both models were trained on a proprietary corpus derived from both public and private sources. They have a context window of 200,000 tokens. Their distinction resides in the fact that Haiku is designed for immediate responses, while Sonnet 3.0 possesses the capability to manage complex tasks thanks to its architectural framework. These models adhered to the Constitutional AI framework, ensuring their alignment with the principles of helpfulness, honesty, and non-harmfulness
GPT-3.5 Turbo (GPT-3.5) and GPT-4 Turbo (GPT-4), both developed by OpenAI, exhibit specific distinctions in their design and capabilities (OpenAI et al., , 2023). GPT-3.5 Turbo operates with a context window of 16,000 tokens, whereas GPT-4 Turbo features a larger context window of 128,000 tokens. Furthermore, GPT-4 Turbo is capable of processing both textual and visual inputs, whereas GPT-3.5 Turbo is restricted to textual data exclusively. Additionally, GPT-4 was trained on a larger and more heterogeneous dataset compared to GPT-3.5 Turbo. Despite this, GPT-3.5 Turbo continues to be a practical and economical choice for numerous applications because of its effective combination of performance and efficiency.
@@ E.2 Decoder-Smal FT
The number of parameters in this category ranges from 1 billion to 3.8 billion. However, the context window and the number of training tokens differ across models. For instance, Olmo-1B-Instruct (Olmo-1B) (Groeneveld et al., , 2024), with 1 billion parameters, and TinyLlama-1.1B-Chat-v1.0 (TinyLlama) (Zhang et al., 2024b, ), with 1.1 billion parameters, share the same context window size of 2,048 tokens and were trained on 3 trillion tokens. In contrast, SmolLM2141414SmolLM2 - https://github.com/huggingface/smollm/blob/main/README.md, which has 1.7 billion parameters, also employs a context window of 2,048 tokens but was trained on a substantially larger dataset of 11 trillion tokens. These models also vary in their training sources: Olmo-1B was trained on subsets of Dolma v1.7 (Soldaini et al., , 2024), TinyLlama utilized the architecture and tokenizer of Llama 2 (Touvron et al., , 2023), and SmolLM2 was trained on a diverse dataset that includes textbooks, web content, code, mathematics, and external data sources.
On the other hand, Llama3.2-3-Instruct (Llama3.2-3B) [^15] Llama3.2-3B-Instruct - https://huggingface.co/meta-llama/Llama-3.2-3B-Instruct was trained on a dataset comprising 15 trillion tokens, with a parameter count of 3 billion and a context window extending up to 8,000 tokens. In contrast, Phi3.5-mini-Instruct (Phi3.5) [^16] Phi3.5-mini-Instruct - https://huggingface.co/microsoft/Phi-3.5-mini-instruct was trained on 3.4 trillion tokens, incorporating 3.8 billion parameters and a significantly larger context window of 128,000 tokens. Gemma2-2B-Instruct (Gemma2-2B) shares the same characteristics as those described in the category decoder OSL, differing only in its data source [^17] Gemma2-2B-Google - https://huggingface.co/google/gemma-2-2b-it. The only merge model in this category is Arcee-Lite [^18] Arcee-Lite - https://huggingface.co/arcee-ai/arcee-lite which have 1.5 billion parameters and was developed using Distilkit [^19] Distilkit - https://github.com/arcee-ai/DistillKit. It supports a context window of 32,000 tokens and its distillation source is Phi-3-Medium (Abdin et al., , 2024).
@@ E.3 Decoder-Medium FT
For merged models, this category includes Arcee-Spark [^21] Arcee-Spark - https://huggingface.co/arcee-ai/Arcee-Spark and Llama-3.1-SuperNova-Lite (SuperNova-Lite) [^22] SuperNova-Lite - https://huggingface.co/arcee-ai/Llama-3.1-SuperNova-Lite. Arcee-Spark, initialized from the Qwen2-7B-Instruct (Yang et al., , 2024), comprises 7 billion parameters. On the other hand, SuperNova-Lite, with 8 billion parameters, is built upon the Llama-3.1-8B-Instruct [^23] Llama-3.1 - https://ai.meta.com/blog/meta-llama-3-1/ architecture and distilled from the Llama-3.1-405B-Instruct model. Both models have a context window of 128,000 tokens.
@@ E.4 Decoder-Large FT
This category includes models with parameters ranging from 9 billion to 14 billion. In this sense, Mistral-Nemo-Instruct-2407 (Mistral-Nemo) [^24] Mistral Nemo - https://mistral.ai/news/mistral-nemo/ has 12 billion parameters and a 128,000 token context window. Its training dataset incorporates a mix of multilingual text, code data, and conversational-style data to ensure high-quality input. Whereas, SuperNova-Medius [^25] Mistral Nemo - https://huggingface.co/arcee-ai/SuperNova-Medius is a merged model of 14 billion parameters based on the Qwen2.5-14B-Instruct architecture. It combines knowledge from both the Qwen2.5-72B-Instruct model and the Llama-3.1-405B-Instruct model through distillation. It has a context windows of 131,072 tokens. In contrast, Gemma2-9-Instruct (Gemma2-9B) has the same features as in Gemma, but differs on its source [^26] Gemma2-9B-Google - https://huggingface.co/google/gemma-2-9b-it and the number of parameters which are 9 billion. The only proprietary model is GPT-4o-mini-2024-07-18 (GPT-4o-mini) [^27] GPT-4o-mini - https://openai.com/index/gpt-4o-mini-advancing-cost-efficient-intelligence/ which is part of the GPT-4o family and is designed to be a cost-efficient, high-performance AI model. It has multimodal capabilities and its context window of 128,000 tokens. It is designed to replace GPT-3.5 Turbo in ChatGPT due to its improved performance and cost-efficiency for various AI applications.